News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 31 October 2008 8.30am EDT

During the last week Chandra completed the observing schedule as planned.

On Oct 28 telemetry indicated an eclipse flag in an unexpected state for 6 seconds. Analysis showed the state change to be the result of unusually low current draw on the spacecraft power bus. This met the threshold for the flight software to set the eclipse flag momentarily. The event had no impact on operations. The current condition is thought to be related to the long-term spacecraft heating trend (and other factors) and the Flight Team will investigate a K-constant flight software patch to prevent future occurrences.

A Chandra press release was issued on 30 Oct describing Chandra observations of the Bullet Cluster that were combined with data from the Compton Gamma-Ray Observatory to search for the signal for primordial antimatter. For further details see: http://chandra.harvard.edu/press/08_releases/press_103008.html

The schedule of targets for the next week is shown below and includes observations of GM Aur coordinated with Gemini North and SS7317 coordinated with Suzaku.

------------------------------------------
Radiation Belts                    Nov  2
Vega                   HRC-I       Nov  3
Vega (3 obs)           HRC-S
31Com                  HRC-I
NGC1232                ACIS-I
MZ10451                ACIS-S      Nov  4
Q0105+061              ACIS-S
Radiation Belts                    Nov  5
SS7317                 ACIS-S/HETG
NGC1232                ACIS-I
NGC4051                ACIS-S/HETG Nov  6
0106+013               ACIS-S
        NA                                 Nov  7
SN2005bx               ACIS-S
Radiation Belts                    Nov  8
M31                    HRC-I
NGC4051                ACIS-S/HETG
0106+013               ACIS-S
GM Aur                 ACIS-S      Nov  9
SDSSJ0148+0001         ACIS-S
NGC4051                ACIS-S/HETG

------------------------------------------

All spacecraft subsystems continued to support nominal operations.

Return to Status Reports